40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 24 22:14:08 2012, Koji, Update, General, AP table cleaning up APTable121124.pdfAPTable121124.graffle.zip
    Reply  Tue Nov 27 03:26:00 2012, Jenne, Update, General, POP beam on POP55, POP camera 
    Reply  Tue Nov 27 17:45:42 2012, Steve, Update, General, spare optics of AP table moved to cabinet IMG_1819.JPG
    Reply  Wed Mar 28 16:44:27 2018, Steve, Update, General, AP table today AP_Table_20180328.png
Message ID: 7752     Entry time: Tue Nov 27 03:26:00 2012     In reply to: 7740
Author: Jenne 
Type: Update 
Category: General 
Subject: POP beam on POP55, POP camera 

Quote:

Some obvious things to be fixed

-  We need POP55 and POP CCD for diagnoses.

 Done.  The beam is also going vaguely in the direction of POP110, but I can't see the beam, so it's tricky. 

Order of operations:

1. Find POP on the table, place iris so I wouldn't forget.  Find beam by putting big IR card where I think beam should be, look at IR card with IR viewer.

2. Move and re-clamp 2" lens so beam is on center of lens.

3. Move and re-clamp 1st (2") mirror so that beam is on center of mirror. 

4. Remove BS-33% so that all the beam goes to POP55, steer that 1st mirror so beam is on POP55's little mirror.  Align little mirror so beam is centered on POP55 (as seen by looking at PD with viewer, finding "edges" of PD, going back to center).

5. Put BS-33% back in place.  The reflected portion of this beam is not possible to see using card+viewer technique.

6. Remove BS-50% that reflects half of this beam to POP110.  Find beam reflected from BS-33% by waving POP camera around.  Steer BS-33 until beam goes back in the direction that the camera used to be mounted.  Adjust camera mount and BS-33 so that beam is on camera.

7. Put BS-50% back in place.  Steer it around with voltmeter on PD to see if beam ever hits PD.  Unsuccessful. Give up, since we have POP55, and POP camera.

8. Make a youtube video: POP, AS, REFL, ITMXF (all on Quad3) - PRMI coarsely aligned, no IFO parts locked.  MICH was locked earlier, but not during video time.

ELOG V3.1.3-